Two spherical conductors A 1 and A 2 of radii r 1 and r 2 are placed concentrically in air. The two are connected by a copper wire as shown in fig. Then the equivalent capacitance of the system is-

Grade:12

1 Answers

Saket trivedi
5 years ago
Once you connect the outer and inner shell, all the charge on inner shell moves to outer shell same as the case that happens with van DE graphs generator.  Inner sphere will have no charge and thus by
C=q/v
V =kq/b
Here b radius of outer shell
Gives:
C=b/k
C= 4